### v4

# Event2Vec: Processing Neuromorphic Events Directly by Representations in Vector Space

Abstract
Neuromorphic event cameras possess superior temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, their asynchronous and sparse data format poses a significant challenge for conventional deep learning methods. Existing methods either convert the events into dense synchronous frame representations for processing by powerful CNNs or Transformers, but lose the asynchronous, sparse and high temporal resolution characteristics of events during the conversion process; or adopt irregular models such as sparse convolution, spiking neural networks, or graph neural networks to process the irregular event representations but fail to take full advantage of GPU acceleration. Inspired by word-to-vector models, we draw an analogy between words and events to introduce event2vec, a novel representation that allows neural networks to process events directly. This approach is fully compatible with the parallel processing capabilities of Transformers. We demonstrate the effectiveness of event2vec on the DVS Gesture, ASL-DVS, and DVS-Lip benchmarks, showing that event2vec is remarkably parameter-efficient, features high throughput and low latency, and achieves high accuracy even with an extremely low number of events or low spatial resolutions. Event2vec introduces a novel paradigm by demonstrating for the first time that sparse, irregular event data can be directly integrated into high-throughput Transformer architectures. This breakthrough resolves the long-standing conflict between maintaining data sparsity and maximizing GPU efficiency, offering a promising balance for real-time, low-latency neuromorphic vision tasks. The code is provided in https://github.com/Intelligent-Computing-Lab-Panda/event2vec.
Neuromorphic Computing, Event Camera
1 Introduction
Neuromorphic computing is an emerging research field that seeks to develop the next generation of artificial intelligence by emulating the brain’s principles (Mead, 1990). A significant advancement stemming from this paradigm is the event camera, a sensor inspired by the biological retina (Gallego et al., 2022). Prominent examples include the Dynamic Vision Sensor (DVS) (Lichtsteiner et al., 2008) and the Asynchronous Time-based Image Sensor (ATIS) (Posch et al., 2011). Unlike traditional cameras that capture synchronous frames, event cameras operate asynchronously, generating events in response to per-pixel brightness changes. This operational principle endows them with exceptionally high temporal resolution (on the order of microseconds), low power consumption, and a High Dynamic Range (HDR) exceeding 120 dB. This asynchronous operation results in a sparse stream of events, typically encoded in the Address-Event Representation (AER) format. An event is represented as a tuple $(x,y,t,p)$, composed of the pixel’s spatial coordinates $(x,y)$, a timestamp $t$, and a binary polarity $p$ that indicates the direction of the brightness change.
Most contemporary deep learning models are designed to operate on dense, regularly structured, multi-dimensional tensors. This regular paradigm is foundational to mainstream deep learning (LeCun et al., 2015) and is ubiquitously employed in modern scientific computing and machine learning frameworks, including NumPy (Harris et al., 2020), TensorFlow (Abadi et al., 2016), and PyTorch (Paszke et al., 2019). Consequently, the sparse and asynchronous nature of event streams in the AER format is fundamentally incompatible with these regular methods. To address this disparity, substantial research efforts have been devoted to converting events to dense representations, or designing new data and network structures to process the irregular events directly.
Figure 1: Conceptual analogy between words and events.
Existing methods primarily address the challenge of event encoding: how to effectively extract information from events and represent it for processing by neural networks. This challenge is analogous to word encoding in natural language processing, a problem successfully addressed by word-to-vector (word2vec) (Mikolov et al., 2013). The word2vec model embeds each word into a fixed-length vector, enabling the relationships between words to be represented by mathematical operations between vectors. This vector representation approach is highly compatible with deep learning architectures and has become a foundational component of modern Natural Language Processing (NLP) models (Devlin et al., 2019a; Brown et al., 2020). As illustrated in Figure 1, we identify numerous parallels between words and events. The key similarities are as follows:
  1. (1)
    Each element is a composite of an index and a position. In NLP, each word is assigned a unique index from a vocabulary, a conversion handled by a tokenizer; the indices in Figure 1, for instance, are generated by the Llama-3 tokenizer (Grattafiori et al., 2024). A word’s position is its sequential location within the sentence (e.g., the word “how” is at position 0 in “how are you”). Similarly, an event’s index is its spatial address, represented by the tuple $(x,y,p)$. Crucially, its position is not the sequence number, but its timestamp $t$, which marks its precise temporal location in the event stream.
  2. (2)
    The set of possible indices is finite. The vocabulary of a language, which forms the dictionary used in NLP, is finite. Likewise, an event camera has a limited set of possible event indices, defined by its sensor’s properties. For example, a DVS128 camera has $2\times 128\times 128$ unique indices, corresponding to 2 polarities across a $128\times 128$ spatial resolution.
  3. (3)
    The sequence exhibits a natural ordering. Words in a sentence are arranged in a specific sequence that dictates meaning. Analogously, events are naturally ordered by their timestamps, reflecting the chronological progression of captured changes. This inherent temporal order is a key characteristic that distinguishes event data from unordered data structures like point clouds.
  4. (4)
    The meaning of an element is determined by its context. A word can be polysemous; for instance, “transformer” can refer to a neural network architecture or a character in an animated series; its specific meaning is disambiguated by the surrounding text. An individual event merely indicates a brightness change at a specific pixel and time, conveying little information in isolation. However, when viewed within a spatiotemporal stream, a sequence of events can delineate an object’s contour, thus giving a single event a higher-level meaning, such as being part of an edge. Therefore, the significance of an event is also fundamentally context-dependent.
Inspired by word2vec, we propose event-to-vector (event2vec), an efficient spatio-temporal representation for asynchronous events. Our contributions are as follows:
  1. (1)
    By embedding events into a vector space, our method natively handles the sparse nature of the input stream, avoiding dense intermediate representations like event frames. This allows for efficient, GPU-accelerated processing with modern network architectures.
  2. (2)
    We propose a parametric spatial embedding and a convolution-based temporal embedding method to capture neighborhood similarity—a characteristic that is critical for accuracy but difficult for a standard embedding layer (a look-up table) to learn.
  3. (3)
    We validated our method on three widely used classification benchmarks: DVS Gesture (Amir et al., 2017), ASL-DVS (Bi et al., 2019), and DVS-Lip (Tan et al., 2022). It achieved competitive accuracy while demonstrating remarkable parameter efficiency, throughput, latency, as well as robustness against a low number of events or low spatial resolutions.
Existing methods either leverage CNNs or Transformers to process dense representations of events, which exploit extremely high computational efficiency on GPUs and enable large-scale model deployment, but sacrifice the sparse, asynchronous, and high temporal resolution characteristics inherent to events; or adopt irregular network architectures (e.g., sparse convolution, spiking neural networks, graph neural networks) to handle raw or downsampled sparse representations of events, yet these irregular models can hardly take full advantage of the massive parallel computing capability of GPUs. Our proposed method, event2vec, addresses this dilemma for the first time, allowing the simultaneous use of sparse irregular representations and dense regular models, and thus provides a brand-new event processing paradigm for neuromorphic vision.
2 Related Work
2.1 Dense Representations and Processing of Events
Dense representations, derived from raw event streams, are fully compatible with conventional deep learning methods. This is typically achieved by integrating events along the time axis to form dense 3D or 4D tensors, such as event frames (Liu and Delbruck, 2018), multi-channel images (Barchid et al., 2022), voxel grids (Bardow et al., 2016), volumetric cubes (Cordone et al., 2022), and patches (Sabater et al., 2023; Peng et al., 2023). Specifically, event-to-frame methods accumulate events within discrete time intervals. The resulting frames can then be processed directly by standard neural networks. However, a significant drawback of these methods is the degradation of the high temporal resolution inherent to event data. This occurs because individual event timestamps are aggregated or quantized during the conversion process. Furthermore, transforming the data into a dense representation negates the inherent spatial sparsity of events. For instance, the generated frames often contain a substantial number of zero-valued pixels. These pixels, while carrying no information, still incur significant memory and computational overhead. While many methods use timestamps implicitly to define the integration interval, some approaches explicitly leverage them to generate temporal weights (Zhu et al., 2019; Gehrig et al., 2019). Finally, the conversion process itself can be computationally intensive, introducing considerable latency that is often prohibitive for real-time applications (Rebecq et al., 2019; Gallego et al., 2022).
2.2 Irregular Representations and Processing of Events
Conversely, methods for processing irregular representations aim to preserve the inherent sparsity and asynchronicity of event data. This category includes Spiking Neural Networks (SNNs) (Maass, 1997; Roy et al., 2019), Sparse Convolutional Networks (Sparse CNNs) (Messikommer et al., 2020; Santambrogio et al., 2024), Graph Neural Networks (GNNs) (Bi et al., 2019; Schaefer et al., 2022), and point-based methods (Yang et al., 2019; Sekikawa et al., 2019; Lin et al., 2023; Ren et al., 2025).
When deployed on neuromorphic hardware (Merolla et al., 2014; Davies et al., 2018), SNNs can process events in a naturally asynchronous, event-driven manner. However, on standard hardware, GPU-based simulations of SNNs produce dense tensor outputs, as the hardware necessitates synchronous processing with discrete time-steps. Consequently, training SNNs on GPUs typically occurs in a synchronous fashion, leading to an unavoidable performance gap between synchronous training and asynchronous inference (Yao et al., 2024; Du et al., 2025). Moreover, the reliance on backpropagation-through-time renders the training process slow and memory-intensive. Sparse CNNs leverage the inherent sparsity of event data, achieving a theoretically low number of Floating-Point Operations (FLOPs). Nevertheless, the architecture of GPUs is not optimized for the dynamic computations and unstructured memory access patterns required for efficient sparse acceleration. Consequently, similar to SNNs, Sparse CNNs fail to fully exploit the massive parallel processing capabilities of GPUs.
Event-based GNNs construct graphs from incoming events, an approach that effectively preserves the spatio-temporal relationships between them. Since empty regions with no event activity do not generate graph nodes, the data’s sparsity is well-utilized. Their main disadvantage lies in the need for careful hyper-parameter tuning, such as the event downsampling rate and neighborhood radius for graph construction. Additionally, functioning as low-pass filters (Nt and Maehara, 2019), GNNs are susceptible to the over-smoothing problem (Zhou et al., 2020), which limits their ability to form deep architectures comparable to modern CNNs and Transformers (Vaswani et al., 2017). Point-based methods treat events from event cameras as analogous to point clouds from Light Detection and Ranging (LiDAR) sensors. A fundamental limitation of most point cloud models is their permutation invariance, which necessitates treating the input as an unordered set. Consequently, the event timestamp is typically relegated to being an additional positional coordinate, thereby discarding the crucial causal ordering of events. To manage the data volume, these methods often employ classic point cloud pre-processing techniques like farthest point sampling, which further increases latency.
3 Methods
3.1 Representing Events in a Vector Space
Leveraging the strong analogy between words and events, we propose a method for representing events within a vector space, which we term event-to-vector (event2vec). An event, generated by a camera with a spatial resolution of $H\times W$, is represented as a tuple $(x,y,t,p)$. For our embedding, we treat the triplet $(x,y,p)$ as the spatial coordinate and the timestamp $t$ as the temporal coordinate. The general formulation for the event2vec embedding is defined as:
v $=\textbf{v}_{s}+\textbf{v}_{t}=\text{Embed}_{s}(x,y,p)+\text{Embed}_{t}(t),$ (1)
where $\textbf{v}\in\mathbb{R}^{D}$ is the resulting $D$-dimensional embedding vector, $\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)\in\mathbb{R}^{D}$ is the spatial embedding vector, and $\textbf{v}_{t}=\text{Embed}_{t}(t)\in\mathbb{R}^{D}$ is the temporal embedding vector. As shown in Eq. 1, this method fuses spatial and temporal information through addition. This additive fusion strategy is directly inspired by the positional encoding mechanism prevalent in Transformers.
3.2 Spatial Embedding
A straightforward approach for the spatial embedding module is to adapt the standard embedding layer from NLP, which is efficiently implemented as a look-up table:
$\textbf{v}_{s}=\text{Embed}_{s}(x,y,p)=\textbf{W}_{s}[p\cdot H\cdot W+y\cdot W+x],$
where $\textbf{W}_{s}\in\mathbb{R}^{(2\cdot H\cdot W)\times D}$ is the learnable embedding matrix and $D$ is the embedding size. This method maps each unique spatial coordinate to a distinct row index in the embedding matrix $\textbf{W}_{s}$.
However, this standard embedding layer imposes no inductive bias on the relationship between indices, compelling the model to learn all spatial relationships from data alone. In a tokenizer, a word’s index is a non-semantic identifier, the assignment of which is primarily determined by the word’s frequency in the training corpus. Consequently, the words at indices $i$ and $i+1$ share no inherent semantic similarity. This assumption does not hold for event coordinates. Images are continuous two-dimensional functions (Gonzalez, 2009). Spatially adjacent pixels are known to exhibit strong correlation. Therefore, an effective spatial embedding should incorporate this locality bias, ensuring that events with close coordinates yield similar embedding vectors:
$\text{Embed}_{s}(x+\Delta x,y+\Delta y,p)-\text{Embed}_{s}(x,y,p)\approx\textbf{0},$
for small coordinate perturbations $(\Delta x,\Delta y)$.
The standard embedding in Eq. 2 fails to account for this crucial spatial relationship, which can impede the learning process. To solve this issue, we propose an elegant parametric algorithm for generating the embedding matrix $\textbf{W}_{\phi}$ via a neural network $\phi$. To systematically enumerate all spatial coordinates within a $P\times H\times W$ volume (where $P=2$ represents the two polarities), we first establish a linear index sequence $\textbf{c}=[0,1,\dots,P\cdot H\cdot W-1]$. This sequence is then decomposed into three probe tensors, $\textbf{x}_{c}$, $\textbf{y}_{c}$, and $\textbf{p}_{c}$, which correspond to the coordinates along the width, height, and polarity dimensions, respectively. The transformation is defined as follows: $\textbf{x}_{c}=\textbf{c}\pmod{W},\textbf{y}_{c}=\left\lfloor\frac{\textbf{c}}{W}\right\rfloor\pmod{H},\textbf{p}_{c}=\left\lfloor\frac{\textbf{c}}{WH}\right\rfloor$. Finally, these probe tensors are passed through $\phi$, which generates the complete embedding matrix $\textbf{W}_{\phi}=\phi(\textbf{x}_{c},\textbf{y}_{c},\textbf{p}_{c})$. By substituting the parametrically generated matrix $\textbf{W}_{\phi}$ into the look-up mechanism of Eq. 2, we establish a direct equivalence for any given event coordinate $(x,y,p)$:
$\textbf{W}_{\phi}[p\cdot H\cdot W+y\cdot W+x]=\phi(x,y,p).$ (4)
Crucially, the parametric network $\phi$ is designed to be a continuous and differentiable function. This property allows us to formally analyze the relationship between neighboring embeddings using a first-order Taylor series expansion:
$\phi(x+\Delta x,y+\Delta y,p)-\phi(x,y,p)$
$\approx\frac{\partial\phi}{\partial x}\Delta x+\frac{\partial\phi}{\partial y}\Delta y+o(\|\Delta\|),$ (5)
where $o(\|\Delta\|)$ represents higher-order remainder terms. As Eq. 5 illustrates, for small perturbations $(\Delta x,\Delta y)$, the difference between the embeddings is approximated by the inner product of the gradient of $\phi$ and the perturbation vector. Consequently, as the perturbations approach zero, this difference vector also approaches zero. In this manner, a continuous parametric network $\phi$ inherently embeds the desired neighborhood semantics, or spatial inductive bias, directly into the embedding matrix. This approach elegantly satisfies the condition outlined in Eq. 3.
3.3 Temporal Embedding
Timestamps, which denote the occurrence time of events, serve a function analogous to positional indices in a sentence. In modern NLP models, relative positional encoding methods (Press et al., 2021; Su et al., 2024) are increasingly favored over absolute methods, such as sinusoidal encoding (Vaswani et al., 2017) or learnable absolute positional embeddings (Devlin et al., 2019b).
However, directly applying these relative positional encoding techniques to event timestamps is ill-suited. Such methods are fundamentally designed for discrete and uniformly spaced indices, whereas event timestamps are continuous and inherently non-uniform. To address this discrepancy, we propose learning the temporal embedding directly from the differences between consecutive timestamps.
Specifically, the temporal embedding module is implemented as a stack of convolutional layers, which take the sequence of the first-order temporal difference of timestamps $\Delta\textbf{t}=[\textbf{t}[1]-\textbf{t}[0],\textbf{t}[2]-\textbf{t}[1],...,\textbf{t}[L-1]-\textbf{t}[L-2],0]$ as the input, where $L$ is the number of events, and the last 0 is the padding value. This design offers several advantages:
  1. (1)
    Time-Shift Invariance: By operating on relative temporal differences, the embedding becomes inherently invariant to absolute shifts in time.
  2. (2)
    Contextual Consistency: The convolutional operations allow the temporal embedding for an event to be influenced by the timing of its immediate neighbors, thereby reinforcing the principle of neighborhood semantics in the time domain. On the other hand, the occurrence of individual events may contain a certain amount of noise, and convolution is applied to achieve the effect of local smoothing and noise reduction.
  3. (3)
    Optimization Efficiency and Inductive Bias: Providing $\Delta\textbf{t}$ as input serves as a form of temporal ‘preconditioning’, aligning with the principle of residual learning (He et al., 2016). While a network could theoretically infer intervals from absolute timestamps t, explicitly modeling $\Delta\textbf{t}$ reduces the optimization burden by providing a direct representation of event velocity. Since the convolution operation essentially performs a weighted summation, it is mathematically congruent with differential inputs $\Delta\textbf{t}$. The summation of consecutive time differences possesses a clear physical interpretation: it represents the accumulated duration of a local event window, enabling the network to directly measure the local event density.
3.4 Event Sampling and Aggregation
Raw event streams often contain an extremely large number of events, with sequence lengths exhibiting substantial variance. Furthermore, deep learning frameworks typically process data in batches, which requires that all tensors within a single batch have uniform dimensions. Consequently, it is necessary to sample or aggregate events from each stream to a fixed-length sequence of size $L$.
In this paper, we primarily use two methods. The first is uniform random sampling. We find that this straightforward method works well in most cases and is extremely computationally efficient. However, a significant limitation of random sampling is the substantial information loss incurred by discarding the majority of the events, leading to suboptimal accuracy in complex tasks. Our second method addresses this by leveraging K-means clustering to aggregate the entire event stream into $L$ representative clusters. Specifically, the clustering process is performed independently on the two event polarities to preserve their distinct information channels. Furthermore, we compute an intensity factor, $\rho$, equal to the number of raw events belonging to that cluster. This intensity factor then modulates the corresponding spatial embedding vector, effectively weighting the representation by its event density.
To reduce the latency of running the K-means cluster algorithm during inference, we propose a GPU-based batched K-means++ algorithm. This method approximates the step-by-step iteration of K-Means++ initialization (Arthur and Vassilvitskii, 2007) via multi-step batch computation. Meanwhile, it leverages the triangle inequality to reduce the computational cost of distance updates, and achieves an efficient GPU-based implementation with PyTorch. Details can be found in Appendix A.1.
3.5 The Formulation of Event2Vec
In summary, the final event2vec representation for a sequence of $L$ events is a tensor $\textbf{V}\in\mathbb{R}^{L\times D}$. The embedding for the $i$-th event in this sequence, $\textbf{V}[i]$, is formulated as:
$\textbf{V}[i]=$ $(\log\bm{\rho}[i]+1)\cdot$
$\bigg(\text{Embed}_{s}(\textbf{x}[i],\textbf{y}[i],\textbf{p}[i])+\text{Embed}_{t}(\Delta\textbf{t})[i]\bigg),$ (6)
where $\bm{\rho}$, x, y, p, and t are sequences representing the intensity factors, spatial coordinates, and timestamps of $L$ events. $\Delta\textbf{t}[i]=\textbf{t}[i+1]-\textbf{t}[i],i=\{0,1,...,L-2\}$ and $\Delta\textbf{t}[L-1]=0$. For a native event, $\bm{\rho}[i]$ is 1, while for a cluster event, it represents the number of raw events aggregated into that cluster. We take the logarithm of $\bm{\rho}$ to suppress clusters with an excessive number of events and prevent them from dominating the entire sequence.
3.6 Network Structure
The proposed methods are validated on widely used event classification tasks with a stacked network including Event2Vec, Backbone and Classification Head. Details can be found in Appendix A.2.
Event2Vec: The spatial embedding module $\phi$ consists of 3 linear layers. It gradually increases the number of features as $3\rightarrow\frac{D}{4}$, $\frac{D}{4}\rightarrow\frac{D}{2}$ and $\frac{D}{2}\rightarrow D$. Layer Normalization (Ba et al., 2016) layers are also inserted after each linear layer to stabilize training. A ReLU activation is placed after the first two Layer Normalizations. The temporal embedding module has a similar structure to the spatial embedding module, except that it replaces linear layers with depth-wise convolutional layers with a kernel size of 3 and a stride of 1, and the numbers of channels gradually increase as $1\rightarrow\frac{D}{4}$, $\frac{D}{4}\rightarrow\frac{D}{2}$ and $\frac{D}{2}\rightarrow D$.
Backbone: Backbone is constructed by stacking multiple Transformer blocks, which consists of a self-attention, a feed-forward network composed of two fully connected layers and an optional pooling to reduce the sequence length. We employ the Forgetting Transformer (Lin et al., 2025) as the self-attention in the backbone. Linear attentions such as Gated Linear Attention (Yang et al., 2024) can also be employed with slight accuracy drop, the results of which are reported in Appendix A.3. It is important to recognize that the forgetting gates in Forgetting Transformer are order-sensitive. To enhance the learning capability, we extend the Forgetting Transformer to a parameter-shared bi-directional formulation. Further details are provided in Appendix A.4.
Classification Head: We employ an average pooling layer to aggregate features across all positions in the sequence. Then a linear layer is used to make a classification decision.
4 Experiments
We conduct a series of experiments on classification tasks using three neuromorphic datasets: DVS Gesture, ASL-DVS, and DVS-Lip. In this section, results are reported in the format $a\pm b$, representing the mean and standard deviation, respectively. For experiments that involve random sampling, results are computed over 10 independent runs on the test set.
4.1 Comparison Between Representations
Accuracy and Parameter Efficiency Table 1 compares the accuracy and model parameters of event2vec with those of other representations across the three datasets. Our models for DVS Gesture and ASL-DVS are trained directly on randomly sampled events. For DVS-Lip, our model first undergoes self-supervised pre-training (refer to Appendix A.5) on cluster events. We then report the fine-tuning accuracy on both randomly sampled events and cluster events generated by the proposed Batched K-means++ algorithm. Our method achieves comparable accuracy on DVS Gesture and the highest accuracy on ASL-DVS and DVS-Lip among other leading representations, while demonstrating exceptional parameter efficiency. For example, previous State-Of-The-Art (SOTA) models use $2.79\times,815.93\times,$ and $12.22\times$ as many parameters as our model on three datasets.
Table 1: Model performance and size comparison on neuromorphic datasets.
Dataset Method + Representation Accuracy (%) Params (MB)
DVS Gesture Sparse GRU + Frame (Subramoney et al., 2023) 97.80 4.80
SNN + Frame (Yao et al., 2023) 98.23 6.50
FARSE-CNN + Window Slicing (Santambrogio et al., 2024) 96.6 10.79
Event MAE + Point Cloud (Sun et al., 2025) 97.75 Unknown
Max-Former + Frame (Fang et al., 2025) 98.6 1.45
[…]
ASL-DVS GNN,CNN + Graph (Bi et al., 2019) 90.10 19.46
GNN & Transformer + Image & Voxel Graph (Yuan et al., 2023) 99.60 220.30
[…]
DVS-Lip ResNet-18 & BiGRU + Frame (Tan et al., 2022) 72.1 241.20
Spiking ResNet18 & BiGRU + Frame (Dampfhoffer and Mesquida, 2024) 75.3 223.63
[…]
(1024 Batched K-Means++ Cluster Events) 75.88
Table 2: Throughput and latency comparison between Event2Vec and previous SOTA models.
Dataset Method Batch Throughput (samples/s) Single-stream Inference
Training Inference Latency (ms) GPU Memory (MB)
Data Pre-processing Forward Total
DVS Gesture Max-Former 241.12 $\pm$   0.55 1077.35 $\pm$       2.20 10.29 $\pm$   0.23 23.89 $\pm$ 11.65 34.18 $\pm$ 11.75 834
Event2Vec 1016.19 $\pm$ 61.18 2900.08 $\pm$   277.30 9.92 $\pm$   5.79 13.51 $\pm$   9.04 23.43 $\pm$ 10.63 602
ASL-DVS GNN & Transformer 78.08 $\pm$   6.39 200.16 $\pm$     12.66 55.12 $\pm$ 38.35 10.90 $\pm$   0.45 66.02 $\pm$ 38.41 5464
Event2Vec 933.58 $\pm$ 16.14 12543.81 $\pm$ 2565.54 1.10 $\pm$   0.23 6.24 $\pm$   2.89 7.34 $\pm$   2.91 824
DVS-Lip Spiking ResNet18 & BiGRU 10.85 $\pm$   0.03 165.29 $\pm$       2.25 373.82 $\pm$   2.21 15.07 $\pm$   7.73 388.89 $\pm$   8.14 1226
Event2Vec 383.71 $\pm$   0.89 942.29 $\pm$     22.36 17.23 $\pm$   3.33 39.87 $\pm$   1.28 57.10 $\pm$   3.62 838
Throughput, Latency and Memory We further compare model throughput and latency between our models and previous SOTA models, and results are shown in Table 2. Appendix A.6 provides more details about these experiments. Throughput is a primary metric governing the efficiency of model training. For different models, we set the batch size to the largest possible power of 2 or the average of two adjacent powers of 2 (e.g., 64, 96, 128, …) without exceeding the GPU memory limit, aiming to maximize the reported throughput. The results demonstrate that event2vec fully leverages the computational efficiency of Transformers, achieving training and inference throughput that is $4.21\times$ and $2.69\times$, $11.96\times$ and $62.67\times$, and $35.36\times$ and $5.70\times$ higher than those of prior works across the three datasets, respectively. For inference tasks on edge devices (e.g., an embedded neuromorphic system), the latency of processing a single event stream and the GPU memory consumed by the model are critical. We conducted comparative experiments and measured three latency components: event data pre-processing latency, model forward propagation latency, and total latency. The results show that across the three datasets, the total latency of our model is only $68\%,11.12\%$, and $14.68\%$ of that of the prior SOTA methods, while the memory consumption is merely $72.18\%,15.08\%$.
4.2 Ablation Experiments
Embedding Comparison We conducted an ablation study on the DVS Gesture dataset to evaluate the accuracy contributions of different components, as detailed in Table 4. We tested various combinations of spatial embedding methods (standard (Eq. 2) vs. parametric (Eq. 4)) and temporal embedding modules (sinusoidal embedding on t vs. convolutional embedding on $\Delta\textbf{t}$). The combination of the standard embedding with our convolutional temporal embedding (Standard + Conv($\Delta\textbf{t}$)) yields the lowest accuracy. We attribute this to the standard embedding layer’s lack of inductive bias, which prevents it from effectively learning neighborhood semantics and subsequently limits the performance of the convolutional temporal encoder. Consequently, when using our parametric embedding, the convolutional encoder achieves the highest accuracy. It is worth noting that our parametric embedding consistently outperforms the standard version when paired with any temporal embedding, validating the effectiveness of incorporating neighborhood semantics.
Robustness to the Number of Events Processing fewer events results in lower resource consumption, which is always desirable in event-based applications. Figure 2 compares our method with the sophisticated sampling techniques from (Araghi et al., 2025), which use a voxel grid representation. The results highlight the inherent effectiveness of event2vec: when paired with simple random sampling, it consistently outperforms the voxel grid representation, even when the latter employs more complex, meticulously designed sampling strategies. Appendix A.7 provides additional details on the variations in the performance of the event2vec model with respect to the number of events.
Figure 2: Accuracy vs. number of events compared to sampling techniques from (Araghi et al., 2025) on the DVS Gesture dataset.
Figure 3: Accuracy vs. spatial resolution: Comparison with the SOTA Max-Former (Fang et al., 2025) on DVS Gesture.
Robustness to Spatial Resolutions We further tested the robustness of event2vec to variations in the spatial resolution of sensors and compared it with Max-Former (Fang et al., 2025) on the DVS Gesture dataset. For event2vec, the coordinates are treated as floating-point numbers, scaled to the target resolution $h\times w$ and then quantized. Subsequently, the coordinates are upscaled back to the original resolution $H\times W$ and re-quantized. For Max-Former, which adopts a frame-based representation, bilinear interpolation is used to downscale the frames to a lower resolution $h\times w$, followed by upscaling back to the original resolution $H\times W$. The results in Figure 3 demonstrate that event2vec exhibits strong robustness to changes in resolution. Moreover, it maintains a classification accuracy significantly higher than random guessing even when the resolution is reduced to $1\times 1$ (i.e., complete loss of spatial information), which indirectly validates the effectiveness of the temporal embedding.
[…]
Table 3: Ablation analysis of embeddings on DVS Gesture.
Spatial Embedding Temporal Embedding Accuracy (%)
[…]
Table 4: K-Means clustering latency and accuracy on DVS-Lip.
Method Latency (ms) Accuracy (%)
[…]
Scikit-learn (iters=300) 383.42 $\pm$ 149.22 75.08
Scikit-learn (iters=20) 374.22 $\pm$ 145.97 74.72
[…]
Cluster Latency Table 4 compares the average per-sample clustering latency on the test set, as well as the test-set accuracy of models trained with clustered data, when different K-Means clustering methods are applied to the DVS-Lip dataset. We compare our approach with two benchmarks: Scikit-learn’s CPU-based K-Means (using K-Means++) and Meta’s Faiss GPU K-Means (Johnson et al., 2019).As shown in Table 4, the proposed batched K-Means++ outperforms benchmarks by achieving peak accuracy with minimal latency. Notably, it is significantly faster than Scikit-learn and more accurate than the GPU-accelerated Faiss. While Faiss (iters=20) offers comparable speed, it suffers a $1.48\%$ accuracy drop compared to our approach.
4.3 Visualization
Neighborhood Semantics To visually inspect the neighborhood semantics, we extract the spatial embedding weights from models trained on the DVS Gesture dataset with the parametric ($\textbf{W}_{\phi}$) and standard ($\textbf{W}_{s}$) embedding layers. For each coordinate $(x,y,p)$, its $D$-dimensional embedding vector is projected onto a 3-dimensional space using Principal Component Analysis (PCA). These 3D vectors are then interpreted as RGB color values and plotted at their corresponding $(x,y)$ locations to form an image. Figure 5(a) visualizes the resulting images for polarity 0 (images for polarity 1 are provided in Appendix A.8). The image derived from $\textbf{W}_{\phi}$ displays smooth, continuous color gradients, akin to a color palette, indicating that spatially adjacent coordinates have semantically similar embeddings. In stark contrast, the image from $\textbf{W}_{s}$ resembles random noise, signifying a lack of learned spatial correlation.
Polarity Similarity An object’s edge moving across a pixel often triggers events of both polarities in close succession. We therefore hypothesize that the embeddings for opposite polarities at the same spatial location should also be semantically related. To test this, we compute the cosine similarity between the embedding vectors of the two polarities at each coordinate. As shown in Figure 5(b), the parametric embedding captures this relationship, exhibiting distinct regions of high similarity. Conversely, the similarity map for the standard embedding is predominantly close to zero, indicating that it fails to learn this inter-polarity correlation.
Vector Field Representation We visualize the learned spatial manifold as a vector field. The $D$-dimensional embedding vectors are projected onto their first two principal components using PCA. These resulting 2D vectors are then visualized using a quiver plot, where each arrow represents the direction and magnitude of the vector at its spatial coordinate. Figure 5(c) illustrates the results. The vector field for the parametric embedding exhibits a coherent, laminar-like flow, revealing a smoothly structured semantic space. In contrast, the field for the standard embedding appears chaotic and turbulent, further confirming its inability to capture meaningful spatial relationships.
Figure 4: Visual comparison of the learned spatial embeddings.
Figure 5: Event-level attention maps on samples from DVS Gesture (Row 1), ASL-DVS (Row 2), and DVS-Lip (Row 3).
Event-wise Attention As event2vec is an event-wise representation, its attention mechanism can be visualized at a fine-grained, event-level resolution. Figure 5 displays attention heatmaps overlaid on the original event streams for DVS Gesture (row 1), ASL-DVS (row 2), and DVS-Lip (row 3). The visualizations reveal that the model correctly focuses on the hands in DVS Gesture, the finger joints and contours in ASL-DVS, and the lip region in DVS-Lip. However, consistent with the lower classification accuracy compared to the other two datasets, we also observe instances where the model incorrectly allocates significant attention to other facial features, such as the eyes and ears.
5 Conclusions
Neuromorphic event cameras introduce a paradigm shift in computer vision, presenting both unique opportunities and significant challenges. A central challenge has been reconciling their asynchronous, sparse nature with the synchronous, dense regular architectures of deep learning. In this paper, we introduced event2vec, a novel representation that directly addresses this challenge by enabling neural networks to natively process asynchronous events. Our experimental results demonstrate that event2vec achieves accuracy competitive with established methods while offering compelling advantages in parameter efficiency, pre-processing overhead, throughput, and robustness across varying numbers of events and spatial resolutions. The remarkable efficiency and robustness of event2vec suggest its significant potential for real-time deployment on resource-constrained edge devices, where low-latency sensing and low-power consumption are paramount. Beyond these performance metrics, the most significant contribution of event2vec is its conceptual alignment of event streams with the paradigm of natural language processing. This opens new avenues for research and application. By treating events as a sequential language, we can begin to explore novel applications by leveraging the sophisticated architectures developed for large language models.
Acknowledgments
This work was supported in part by CoCoSys, a JUMP2.0 center sponsored by DARPA and SRC, the National Science Foundation (CAREER Award, Grant #2312366, Grant #2318152), the DARPA Young Faculty Award and the DoE MMICC center SEA-CROGS (Award #DE-SC0023198).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Batched K-Means++ Cluster Algorithm
For challenging classification tasks like DVS-Lip, random sampling leads to significant information loss, resulting in suboptimal accuracy. To ensure that all events contribute to the final representation, we employ an event clustering approach. Given a raw event stream $\mathcal{E}=\{(\mathbf{x},\mathbf{y},\mathbf{t},\mathbf{p})\}$ containing $N$ events, our objective is to generate $L$ cluster event streams $\mathcal{R}=\{(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{c},\bm{\rho})\}$, where $(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{c})$ denotes the coordinates of the cluster centers and $\bm{\rho}$ represents the event count within each cluster. Notably, we perform clustering separately for the two polarities to prevent the loss of physical significance caused by polarity mixing. Specifically, assuming the number of events for each polarity is $N_{0}$ and $N_{1}$, we calculate the numbers of clusters as $L_{0}=\text{Round}(\frac{N_{0}}{N}\cdot L)$ and $L_{1}=L-L_{0}$, respectively. Finally, the two resulting cluster streams are merged and sorted chronologically.
The general practice of K-Means clustering involves using the K-Means function provided in Scikit-learn (sklearn)(Pedregosa et al., 2011), a Python machine learning library. However, this function is implemented on the CPU, resulting in slow execution when the number of events is large, which significantly increases the latency of the model in processing real-time tasks. To address this issue, we propose a GPU-based Batched K-Means++ Event Clustering algorithm, whose detailed workflow is presented in Algorithm 1. The acceleration of this algorithm mainly stems from the following optimizations:
  1. 1.
    Reduced Loop Overhead: The number of loop iterations is reduced by a factor of $B$.
  2. 2.
    Parallel Computing: Utilizes torch.cdist to compute distances from all points to the batch of $B$ new centers in parallel.
  3. 3.
    Incremental Update: The update of $\mathbf{D}^{2}$ leverages a variant of the triangle inequality, requiring comparison only between the current known minimum distance and the distance to the newly added batch of centers, avoiding recomputation of all pairwise distances.
1
Input: Raw Event Stream $\mathcal{E}=\{(\mathbf{x},\mathbf{y},\mathbf{t},\mathbf{p})\}$, Total Points NN
Params: Target Cluster Count $L$, Spatial Dimensions $H,W$, Batch Size $B$, Max Iterations $I_{max}$
Output: Downsampled Event Set $\mathcal{S}$
2
/* 1. Data Preprocessing & Normalization */
3 Move the data to the GPU
4 Compute normalized time $\hat{\mathbf{t}}=\frac{\mathbf{t}-t[0]}{t[N-1]-t[0]}$
5 Construct 3D feature space point set $\mathbf{V}=\{(\frac{x[i]}{W},\frac{y[i]}{H},\hat{t}[i])\}_{i=0}^{N-1}$
6 Split $\mathbf{V}$ into positive set $\mathbf{V}^{0}$ and negative set $\mathbf{V}^{1}$ based on polarity $\mathbf{p}$
7 Allocate target center counts $L_{0}$ and $L_{1}$
8 Initialize result set $\mathcal{R}=\emptyset$
9
/* Cluster for each polarity separately */
10 for each point set $\mathbf{V}_{sub}\in\{\mathbf{V}^{0},\mathbf{V}^{1}\}$ and target count $K_{sub}\in\{L_{0},L_{1}\}$ do
11 if $|\mathbf{V}_{sub}|==0$ then
12    continue
13   end if
14
 /* Phase 1: Batched K-Means++ Initialization */
15   Randomly select the first center $\mathbf{c}_{0}\in\mathbf{V}_{sub}$, initialize center set $\mathbf{C}=\{\mathbf{c}_{0}\}$
16   Compute squared distance from all points to first center $\mathbf{D}^{2}=\|\mathbf{V}_{sub}-\mathbf{c}_{0}\|^{2}$
17
18 while $|\mathbf{C}|<K_{sub}$ do
19      Calculate sample count for this batch $M=\min(B,K_{sub}-|\mathbf{C}|)$
    /* Parallel Sampling: Use current distance as probability weights */
20      Sample $M$ new candidate centers $\mathbf{C}_{new}$ based on weights $\mathbf{w}\propto\mathbf{D}^{2}$
21      Add $\mathbf{C}_{new}$ to $\mathbf{C}$
    /* Incremental Distance Update: Only to newly added centers */
22    $\mathbf{D}^{2}_{new}=\min_{c\in\mathbf{C}_{new}}\|\mathbf{V}_{sub}-c\|^{2}$
23      Update global minimum distance $\mathbf{D}^{2}\leftarrow\min(\mathbf{D}^{2},\mathbf{D}^{2}_{new})$
24
25   end while
26
 /* Phase 2: Standard Lloyd’s Iteration */
27 for $iter=0$ to $I_{max}-1$ do
28    E-step: Assign labels to points based on the nearest center in $\mathbf{C}$
29    M-step: Compute centroids of each cluster as the new $\mathbf{C}$
30    if the center shift <tol<tol then
31       break
32      end if
33
34   end for
35
 /* Denormalization & Intensity Calculation */
36   Count points in each cluster as Intensity $\bm{\rho}$
37   Map coordinates of $\mathbf{C}$ back to physical dimensions $(W,H,t_{span})$
38   Add result $(\mathbf{x}_{c},\mathbf{y}_{c},\mathbf{t}_{c},\mathbf{p}_{sub},\bm{\rho})$ to $\mathcal{R}$
39
40 end for
41
/* 3. Post-processing */
42 Merge all results in $\mathcal{R}$
43 Sort results by time $\mathbf{t}_{c}$ to ensure memory contiguity
44 Return Sorted Tensor Data
Algorithm 1 Batched K-Means++ Event Clustering on GPU
A.2 Model Structures and Hyper-parameters
Unless otherwise stated, all models were trained using BFloat16 mixed precision. The training configuration for all models includes a base learning rate of $lr_{b}=0.001$, a batch size of 64, and the AdamW optimizer (Loshchilov and Hutter, 2019) for 64 epochs. The effective learning rate is determined by a linear scaling rule based on the number of GPUs ($n_{gpus}$) used in distributed data-parallel training: $lr=lr_{b}\cdot n_{gpus}/256$. A warmup phase is implemented for the first 4 epochs, during which the learning rate is linearly increased from $0.01\cdot lr$ to $lr$. For the subsequent epochs, a cosine annealing schedule (Loshchilov and Hutter, 2017) is employed to gradually reduce the learning rate to a minimum value, $lr_{min}$. For the DVS Gesture and ASL-DVS datasets, both weight decay and label smoothing were disabled. In contrast, for the DVS-Lip classification task, we set the weight decay to 0.05 and applied label smoothing with a factor of 0.1.
Figure 6 illustrates the detailed network architecture, including (a) the spatial embedding architecture of event2vec, (b) the temporal embedding architecture of event2vec, and (c) the architecture of the entire network. Table 5 provides a detailed summary of the model-specific hyper-parameters. Here, $D$ denotes the embedding dimension, $l$ is the number of Transformer blocks in Backbone, $D_{f}$ represents the hidden feature dimension of the feed-forward neural network (FFN), and $n_{head}$ is the total number of attention heads. The repeats parameter specifies how many times the training set is iterated through within a single epoch. Notably, the number of heads for the key (k) and value (v) projections is set to $n_{head}/2$, and group normalization (Wu and He, 2018) is applied to both. To prevent exploding gradients, we employ gradient clipping, capping the $L_{2}$ norm of the gradients at 1.0.
For the DVS Gesture classification model, the output of each FFN is average-pooled with a window size of 2, whereas other models do not use pooling. The model for the DVS-Lip classification task was pre-trained on the DVS-Lip dataset using a self-supervised learning approach. This pre-training phase utilized a minimum learning rate of $lr_{min}=10^{-6}$, a weight decay of 0.05, a repeats value of 3, and a masking ratio of 30%. Refer to Appendix A.5 for more details.
Figure 6: The network architecture for event classification using the event2vec representation.
Table 5: Hyper-parameters of training models for classification tasks on different datasets.
Dataset $D$ $D_{f}$ $n_{head}$ $l$ Repeats $n_{gpus}$ $lr_{min}$
DVS Gesture 64 128 2 4 24 4 0
ASL-DVS 64 128 2 2 1 7 $10^{-6}$
DVS-Lip 192 384 6 16 3 4 $10^{-6}$
A.3 Accuracy with Different Types of Self-Attention
In addition to the Forgetting Transformer (FoX), we also evaluated the performance when using Gated Linear Attention (GLA), with the results presented in Table 6. The results show that GLA achieves high performance on relatively easy classification tasks such as DVS Gesture and ASL-DVS; however, its performance drops by 3.53% on the more challenging DVS-Lip classification task, which may be attributed to the fact that linear attention struggles to prevent the decay of long-term memory when processing long input sequences. Overall, the performance degradation across all three tasks is negligible when replacing FoX with GLA, indicating that Event2Vec is not sensitive to the specific type of attention mechanism adopted.
Table 6: Comparison of accuracy with different types of self-attentions.
Dataset Accuracy of FoX (%) Accuracy of GLA (%)
[…]
DVS-Lip 75.88 72.35
A.4 Bi-directional Self-Attentions
We evaluated two variants of Self-Attention in our experiments, namely Gated Linear Attention (GLA) (Yang et al., 2024) and Forgetting Transformer (FoX) (Lin et al., 2025). GLA is a typical linear attention mechanism, which can be regarded as a special case of Recurrent Neural Networks (RNNs) (Katharopoulos et al., 2020), where the input sequence order affects the output results. Although FoX does not fall into the category of linear attention, it adopts an RNN-style gating mechanism that depends on input sequence order, thus the input order also exerts an impact on its outputs. Due to the fixed and limited size of hidden states, RNNs inevitably suffer from long-distance information attenuation when processing ultra-long sequences. To mitigate the degradation of long-term memory, we extend GLA and FoX to bidirectional variants.
We adapt this formulation to be bi-directional by inputting both forward and reversed $\textbf{Q},\textbf{K},\textbf{V}$. The bi-directional outputs are computed as:
$\textbf{O}_{f}$ $=\text{Attention}(\textbf{Q},\textbf{K},\textbf{V}),$ (7)
$\textbf{O}_{b}$ $=\text{Reverse}(\text{Attention}(\overleftarrow{\textbf{Q}},\overleftarrow{\textbf{K}},\overleftarrow{\textbf{V}})),$ (8)
$\textbf{O}_{fb}[t]$ $=\textbf{W}_{fb}[\textbf{O}_{f}[t];\textbf{O}_{b}[t]].$ (9)
Unlike classic bi-directional RNNs (Schuster and Paliwal, 1997) that often use independent parameters for each direction, our model employs shared parameters for the two directions. By sharing the projection weights ($\textbf{W}_{q},\textbf{W}_{k},\textbf{W}_{v},\textbf{W}_{forget}$) across both passes, we ensure that the parameter count remains comparable to the uni-directional baseline, with the only increase arising from the fused output projection $\textbf{W}_{fb}$.
We also tested the performance changes when using bi-directional self-attentions without parameter sharing, and the results are summarized in Table 7. The results show that the number of parameters increases by approximately 25% when parameters are not shared. Although the fitting capacity is theoretically improved without parameter sharing, the test-set accuracy instead decreases, indicating slight overfitting. This experimental result demonstrates that our bi-directional self-attentions with parameter sharing not only reduce the number of parameters but also mitigate overfitting.
Table 7: Changes in accuracy and parameters when using bi-directional self-attentions without parameter sharing.
Dataset Parameters (MB) Accuracy (%)
DVS Gesture 0.65 (+25%) 96.63 (-0.94)
ASL-DVS 0.34 (+26%) 99.86 (-0.05)
DVS-Lip 22.90 (+25%) 75.36 (-0.52)
A.5 Self-supervised Training Details
The event-wise nature of the event2vec representation lends itself well to self-supervised pre-training, which can significantly enhance model performance. Specifically, we adopt a masked modeling approach, akin to that used in BERT. The training objective is to mask the spatial coordinates $(x,y,p)$ of a subset of these events and train the model to predict the masked coordinates based on the context provided by the surrounding events and their associated temporal information. This task compels the model to learn a meaningful understanding of spatio-temporal event patterns.
The self-supervised training framework is analogous to the Masked Language Model (MLM) objective in BERT (Devlin et al., 2019a). Given a batch of embedding tensors v of shape $(B,L,D)$, where $B$ is the batch size, $L$ is the sequence length, and $D$ is the embedding dimension, the process begins by randomly masking a portion of the input tokens.
A binary mask m of shape $(B,L)$ is generated from a Bernoulli distribution. The probability of masking any given token is set to $30\%$, which defines the mask ratio. To prevent the model from making predictions via simple interpolation, we mask out $l_{mask}$ consecutive tokens in sequence. The length of each masked span $l_{mask}$ is sampled from a geometric distribution $l_{mask}\sim\text{Geometric}(p)$ with $p=0.1$, resulting in an average length of $10$. Each token $\textbf{v}[i][j]$ corresponding to a mask entry $\textbf{m}[i][j]=1$ is replaced by a single, learnable, $D$-dimensional mask token $\textbf{v}_{m}$. This operation results in a corrupted embedding tensor, denoted as $\hat{\textbf{v}}$. Concurrently, the original coordinates $(\textbf{x}_{m},\textbf{y}_{m},\textbf{p}_{m})$ of the masked tokens are preserved to serve as the ground truth for the reconstruction loss.
The corrupted tensor $\hat{\textbf{v}}$ is then processed by the model’s linear attention layers. Following this, the output embeddings that correspond to the initially masked positions, denoted as $\hat{\textbf{v}}_{m}$ are extracted from the final output tensor using the mask m.
The objective is for the model to reconstruct the original spatial and polarity information from these corrupted embeddings. To achieve this, we first apply the inverse of the spatio-temporal fusion operation to isolate the spatial component of the reconstructed embeddings:
$\hat{\textbf{v}}_{s}=\frac{\hat{\textbf{v}}_{m}}{\bm{\rho}}-\textbf{v}_{t}.$ (10)
[…]
A.6 Experimental Environment for Performance Testing
All experiments related to performance testing (such as throughput and latency) mentioned in this paper were conducted on a Red Hat Enterprise Linux 8.10 server. This server was equipped with an NVIDIA A100 GPU (80GB PCIe), an Intel Xeon Gold 6326 CPU (utilizing 8 cores), and 256GB of RAM. To mitigate the impact of data I/O, all datasets were loaded entirely into RAM for the duration of the experiments.
The three models included for comparison in Table 2 are Max-Former (Fang et al., 2025), GNN & Transformer (Yuan et al., 2023), and Spiking ResNet18 & BiGRU (Dampfhoffer and Mesquida, 2024). All of these models provide official open-source code, which allowed us to conduct experiments based on their codebases.
A.7 Impact of the number of events on DVS Gesture
As illustrated in Figure 7, we benchmark the impact of varying the number of randomly sampled events ($L$) on several key metrics: training/inference throughput, single event stream inference latency, and accuracy on the DVS Gesture dataset. Experimental results on the ASL-DVS dataset show similar trends, and thus are not presented here. As $L$ increases, both training and inference throughput drop sharply, while the single-stream latency remains nearly unchanged. This indicates that for a single sample, the dominant latency arises from the CUDA kernel launch overhead rather than computation itself. Meanwhile, the accuracy improves rapidly with the growth of $L$, demonstrating that more events facilitate the model’s decision-making process.
Figure 7: Effect of number of events on DVS Gesture.
A.8 Visualization of Neighborhood Semantics
Due to space constraints in the main paper, Figures 5(a) and 5(c) display visualizations for only a single event polarity. For completeness, this section provides supplementary visualizations that include both polarities. Figure 8 illustrates the embedding weights mapped to the RGB color space, while Figure 9 depicts them as a vector field.
Figure 8: Visualization of the parametric embedding weight $\textbf{W}_{\phi}$ and the standard embedding weight $\textbf{W}_{s}$ in the RGB domain.
Figure 9: Visualization of the parametric embedding weights $\textbf{W}_{\phi}$ and the standard embedding weight $\textbf{W}_{s}$ as a vector field.
A.9 Data Augmentations
Transformers (including linear attention) lack inductive bias and thus require more data for learning. We used data augmentation methods to expand the data volume to a certain extent, thereby improving performance. Specifically, we did not use data augmentation on the ASL-DVS dataset because we found that SOTA (State-of-the-Art) performance could be achieved without it—this is likely due to the sufficient scale of this dataset: the number of samples in its training set is approximately 80,640, while that of DVS Gesture is 1,176, and DVS-Lip is 14,896.
Denote $\mathcal{U}(a,b)$ as the uniform distribution between $a$ and $b$, and $\text{RandInt}(m,n)$ as a random integer taken from the set $\{m,m+1,...,n\}$, where each integer has an equal probability of being selected.
For an event stream, the data augmentations are applied on events directly. For simplicity, we omit the event index in this subsection. Unless otherwise specified, augmentations are applied on each event stream independently. Note that the coordinates are converted to floating-point percision before applying any augmentation. After all augmentations are applied, only events whose coordinates are valid, i.e., $x\in[0,W-1],y\in[0,H-1]$, are kept.
For the classification task on DVS Gesture, the following transformations are each applied independently with a probability of $0.6$:
  • Random Resizing: Coordinates $(x,y)$ are scaled to $(s_{x}\cdot x,s_{y}\cdot y)$, with scaling factors $s_{x},s_{y}\sim\mathcal{U}(0.8,1.2)$.
  • Random Rotation: Coordinates are rotated by an angle $r\sim\mathcal{U}(-10,10)$ degrees.
  • Random Shearing: A shear transformation is applied with factors $\lambda_{x},\lambda_{y}\sim\mathcal{U}(-0.02,0.02)$.
  • Random Translation: Coordinates are translated by offsets $d_{x},d_{y}\sim\mathcal{U}(-16,16)$.
  • Random Erasing: Erase an $h\times w$ area with $h,w\sim\mathcal{U}(0,16)$ with a probability of 0.1. The center of this area $(c_{x},c_{y})$ satisfy $c_{x}\sim\mathcal{U}(0,W-1),c_{y}\sim\mathcal{U}(0,H-1)$.
  • Temporal Chunk Dropout: A number of temporal chunks, $n_{r}=\text{RandInt}(0,8)$, are removed from the event stream. The length of each removed chunk, $l_{chunk}$, is determined relative to the total stream length, $L$, according to the sampling distribution $l_{chunk}=\frac{\text{RandInt}(1,256)}{L}$.
During the self-supervised phase of the model for classifying DVS-Lip, a series of geometric transformations are employed. Each of the following augmentations is applied independently with a probability of $0.5$:
  • Random Resizing: Coordinates $(x,y)$ are scaled to $(s_{x}\cdot x,s_{y}\cdot y)$, with scaling factors $s_{x},s_{y}\sim\mathcal{U}(0.8,1.2)$.
  • Random Rotation: Coordinates are rotated by an angle $r\sim\mathcal{U}(-15,15)$ degrees.
  • Random Shearing: A shear transformation is applied with factors $\lambda_{x},\lambda_{y}\sim\mathcal{U}(-0.05,0.05)$.
  • Horizontal Flipping: The event stream is flipped horizontally.
  • Random Translation: Coordinates are translated by offsets $d_{x},d_{y}\sim\mathcal{U}(-16,16)$.
When training the model for classifying DVS-Lip, we use the following data augmentations:
  • Random Resizing: Coordinates $(x,y)$ are scaled to $(s_{x}\cdot x,s_{y}\cdot y)$, with scaling factors $s_{x},s_{y}\sim\mathcal{U}(0.8,1.2)$.
  • Random Rotation: Coordinates are rotated by an angle $r\sim\mathcal{U}(-15,15)$ degrees.
  • Random Shearing: Shear transform on $x$ and $y$ with shear factors $\lambda_{x},\lambda_{y}\sim\mathcal{U}(-0.05,0.05)$.
  • Random Flip: The event stream is flipped horizontally with a probability of $1$.
  • Random Translation: Translate $x$ and $y$ with translations $d_{x},d_{y}\sim\mathcal{U}(-16,16)$.
  • Random Erasing: Erase an $h\times w$ area with $h,w\sim\mathcal{U}(0,16)$ with the probability $0.1$. The center of this area $(c_{x},c_{y})$ satisfies $c_{x}\sim\mathcal{U}(0,W-1),c_{y}\sim\mathcal{U}(0,H-1)$.
  • Random Chunk Drop: Randomly mask $n$ temporal chunks with $n=4$. The length of each chunk $l$ is sampled from $\mathcal{U}(1,128)$ and scaled by the ratio of valid events. The starting position $s$ satisfies $s\sim\mathcal{U}(0,L-1)$.
The augmentations listed above are each applied independently with a probability of $0.5$. The token-mix is applied on the embedding tensor with a probability of $0.5$. Specifically, when training on cluster events, the intensity $\rho$ is randomly set to $1$ with a probability of $0.1$. We use drop path (Larsson et al., 2016) in the linear attention layer, with the probability increasing linearly from 0 to $0.4$ with depth.
In addition, all other methods in Tab 1 also use data augmentation, which is summarized in Tab 8.
Table 8: Data augmentation strategies for methods in Table 1.
Dataset Method + Representation Data Augmentations
DVS Gesture Sparse GRU + Frame (Subramoney et al., 2023) Random crop, translation, and rotation
SNN + Frame (Yao et al., 2023) Random slice and integrate
FARSE-CNN + Window Slicing (Santambrogio et al., 2024) Random coordinate translations
Event MAE + Point Cloud (Sun et al., 2025) Point resampling from Point-BERT (Yu et al., 2022)
Max-Former + Frame (Fang et al., 2025) Mixup and Cutmix
Linear Attention + Event2Vec Random resize, rotation, shear, translate, erase, and chunk dropout
ASL-DVS GNN,CNN + Graph (Bi et al., 2019) Random scale, flip, and rotation of node positions
GNN & Transformer + Image & Voxel Graph (Yuan et al., 2023) Random scale and translate
Linear Attention + Event2Vec None
DVS-Lip ResNet-18,BiGRU + Frame (Tan et al., 2022) Random crop and horizontal flip
Spiking ResNet18 & BiGRU + Frame (Dampfhoffer and Mesquida, 2024) Random crop, horizontal flip, spatial masking, zoom, and temporal mask
Linear Attention + Event2Vec Random resize, rotate, shear, flip, translate, and erase